# KernelSHAP-IQ: Weighted Least Square Optimization for Shapley Interactions

### v1

@@ C.1.2 Language Model (LM)
Language models (LMs) are widely investigated with Shapley interactions (Fumagalli et al., 2023; Kolpaczki et al., 2024; Tsai et al., 2023; Sundararajan et al., 2020) highlighting that explanations based on interactions are more expressive than word-level explanations. Similar to the related work, we also investigate KernelSHAP-IQ’s approximation quality in a LM scenario. The LM used in our experiments is a pre-trained DistilBert (Sanh et al., 2019) sentiment analysis model. The LM used is available at https://huggingface.co/lvwerra/distilbert-imdb via the transformers API (Wolf et al., 2020). The transformer was pre-trained on the IMDB movie review dataset (Maas et al., 2011; Lhoest et al., 2021). Provided a tokenized text input, the model predicts the sentiment of the input with a sentiment score ranging from negative $-1$ to positive $1$. For explanation purposes, we remove features (tokens) from the tokenized representation of the input.
@@ C.1.3 Convolutional Neural Network (CNN)
Similar to Fumagalli et al. (2023); Kolpaczki et al. (2024), we evaluate the approximation quality of KernelSHAP-IQ and the available baselines on a convolutional neural network (CNN) image-input explanation task. The CNN is a ResNet18 (He et al., 2016) fitted on the ImageNet (Deng et al., 2009) dataset. The CNN is available at https://pytorch.org/vision/main/models/generated/torchvision.models.resnet18.html via pytorch (Paszke et al., 2019). The task is to explain the predicted label’s probability for random ImageNet images. Individual pixels are grouped together as superpixels with SLIC (Achanta et al., 2012). Missing features (superpixels) are set to the gray as a method of mean-imputation.
@@ Appendix D Additional Empirical Results
D.1 Runtime Analysis.
Next to the experiments regarding approximation quality, we also conducted a small runtime analysis of KernelSHAP-IQ and all available baseline algorithms. The results are summarized in Figure 5. We run KernelSHAP-IQ, SHAP-IQ, SVARM-IQ, and permutation sampling on the LM and an example sentence consisting of 14 words. We let the approximation algorithms estimate the SII values of order $l=2$. The main computational cost results from the model evaluations (accesses to $\nu$), which is bounded by the model’s inference time. From Figure 5 it is apparent that with increasing the amount of allowed model accesses, the runtime scales linearly for all approximation algorithms and all estimators perform equally.

### v2

@@ C.1.2 Language Model (LM)
Language models (LMs) are widely investigated with Shapley interactions (Fumagalli et al., 2023; Kolpaczki et al., 2024b; Tsai et al., 2023; Sundararajan et al., 2020) highlighting that explanations based on interactions are more expressive than word-level explanations. Similar to the related work, we also investigate KernelSHAP-IQ’s approximation quality in a LM scenario. The LM used in our experiments is a pre-trained DistilBert (Sanh et al., 2019) sentiment analysis model. The LM used is available at https://huggingface.co/lvwerra/distilbert-imdb via the transformers API (Wolf et al., 2020). The transformer was pre-trained on the IMDB movie review dataset (Maas et al., 2011; Lhoest et al., 2021). Provided a tokenized text input, the model predicts the sentiment of the input with a sentiment score ranging from negative $-1$ to positive $1$. For explanation purposes, we remove features (tokens) from the tokenized representation of the input.
@@ C.1.3 Convolutional Neural Network (CNN)
Similar to Fumagalli et al. (2023); Kolpaczki et al. (2024b), we evaluate the approximation quality of KernelSHAP-IQ and the available baselines on a convolutional neural network (CNN) image-input explanation task. The CNN is a ResNet18 (He et al., 2016) fitted on the ImageNet (Deng et al., 2009) dataset. The CNN is available at https://pytorch.org/vision/main/models/generated/torchvision.models.resnet18.html via pytorch (Paszke et al., 2019). The task is to explain the predicted label’s probability for random ImageNet images. Individual pixels are grouped together as superpixels with SLIC (Achanta et al., 2012). Missing features (superpixels) are set to the gray as a method of mean-imputation.
@@ Appendix D Additional Empirical Results
D.1 Runtime Analysis
Next to the experiments regarding approximation quality, we also conducted a small runtime analysis of KernelSHAP-IQ and all available baseline algorithms. The results are summarized in Figure 5. We run KernelSHAP-IQ, SHAP-IQ, SVARM-IQ, and permutation sampling on the LM and an example sentence consisting of 14 words. We let the approximation algorithms estimate the SII values of order $l=2$. The main computational cost results from the model evaluations (accesses to $\nu$), which is bounded by the model’s inference time. From Figure 5 it is apparent that with increasing the amount of allowed model accesses, the runtime scales linearly for all approximation algorithms and all estimators perform equally. Further, we evaluate the runtime of the different approximation methods jointly with their performance. Table 3 shows the mean (averaged over 10 runs) runtime of each approximator to reach certain MSE levels. For CH, KernelSHAP-IQ and Inconistent KernelSHAP-IQ are most efficient in terms of model calls to reach the desired MSE error levels. The runtime is similar for the other baselines. While Inconsistent KernelSHAP-IQ fails to reach the desired error levels for LM, KernelSHAP-IQ performs equally well with SVARM-IQ for LM. Permutation sampling generally has a fast runtime, but does not achieve good estimation qualities in sensible time.
@@ D.1 Runtime Analysis
Table 3: Mean runtime of each approximator to reach certain MSE error levels in terms of model calls and elapsed time in seconds. The runtime is averaged over 10 independent runs for two benchmarks. For LM, Inconsistent KernelSHAP-IQ never reaches the MSE error levels and Permutation requires more than $2^{14}$ evaluations.
Benchmark Error Metric KernelSHAP-IQ Inc. KernelSHAP-IQ SVARM-IQ SHAP-IQ Permutation
[…]
Time (s) 0.023 0.012 0.035 0.033 0.024
[…]
Time (s) 0.026 0.017 0.043 0.037 0.071
[…]
Time (s) 35.5 - 35.6 95.9 -
[…]
Time (s) 70.2 - 71.6 126.1 -